# Zhengdao Li1∗ Siheng Wang2∗ Zeyu Zhang1∗† Hao Tang1‡

### v1

@@ Introduction
Given a training mini-batch =$\mathcal{B}=\{(m_{i},t_{i})\}_{i=1}^{N_{b}}\subseteq\mathcal{D}$ ($N_{b}=|\mathcal{B}|\ll N_{q}$), we compute the momentum features for each sample pair and enqueue them into their respective queues, while simultaneously dequeuing the oldest $N_{b}$ entries to maintain a fixed queue size. For contrastive learning, each motion sample takes its paired text as the positive example, and all entries in $Q^{t}$ are treated as negatives. The motion-to-text contrastive loss is formulated as:
@@
7:  Sample a mini-batch =$\mathcal{B}=\{(m_{i},t_{i})\}_{i=1}^{N_{b}}\subseteq\mathcal{D}$
@@
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the concatenation of both terms, $P$ is the 2D Position embedding, and Q, K, V are the query, key, and value matrices. We refer to this attention mechanism as Semantics Spatial-Temporal Attention (SSTA). A simplified variant that only concatenates the text embedding $t$ with the motion vector $z$ and excludes retrieval features, is termed Spatial-Temporal Attention (STA). As shown in Figure 2(b), both attention mechanisms will be used in the subsequent motion generation module.
@@
Given a motion $\mathbf{m}\in\mathbb{R}^{T\times J}$, we first extract 2D latent features $\hat{\mathbf{y}}\in\mathbb{R}^{T\times J}$ using a 2D convolutional encoder $E_{2d}$. We then apply residual vector quantization (RVQ) (Guo et al. 2024) with $L{+}1$ levels:
@@
[…]
@@
We design our 2D retrieval-augmented masked transformer to generate the base-layer motion tokens $\mathbf{y}^{0}\in\mathbb{R}^{T\times J}$, conditioned on the text embedding $t$, retrieved text feature $R_{t}$, and retrieved motion feature $R_{m}$, all of which are fused using SSTA. To train the model, we randomly mask a subset of tokens in $\mathbf{y}^{0}$, replacing them with a $[\mathrm{MASK}]$ token to obtain a corrupted sequence $\mathbf{y}^{0}_{\text{msk}}$. We perform a 2D Mask strategy (Yuan et al. 2024) by first randomly masking along the temporal dimension and then randomly masking along the spatial dimension on those unmasked frames. The model is then trained to reconstruct the original tokens by minimizing the negative log-likelihood:
@@
[…]

### v2

@@ Introduction
Given a training mini-batch $\mathcal{B}=\{(m_{i},t_{i})\}_{i=1}^{N_{b}}\subseteq\mathcal{D}$ ($N_{b}=|\mathcal{B}|\ll N_{q}$), we compute the momentum features for each sample pair and enqueue them into their respective queues, while simultaneously dequeuing the oldest $N_{b}$ entries to maintain a fixed queue size. For contrastive learning, each motion sample takes its paired text as the positive example, and all entries in $Q^{t}$ are treated as negatives. The motion-to-text contrastive loss is formulated as:
@@
7:  Sample a mini-batch $\mathcal{B}=\{(m_{i},t_{i})\}_{i=1}^{N_{b}}\subseteq\mathcal{D}$
@@
where $\cdot$$\cdot$] denotes the concatenation of both terms, $P$ is the 2D Position embedding, and Q, K, V are the query, key, and value matrices. We refer to this attention mechanism as Semantics Spatial-Temporal Attention (SSTA). A simplified variant that only concatenates the text embedding $t$ with the motion vector $z$ and excludes retrieval features, is termed Spatial-Temporal Attention (STA). As shown in Figure 2(b), both attention mechanisms will be used in the subsequent motion generation module.
@@
Given a motion $\mathbf{m}\in\mathbb{R}^{T\times J}$, we first extract 2D latent features $\hat{\mathbf{y}}\in\mathbb{R}^{T\times J}$ using a 2D convolutional encoder $E_{2d}$. We then apply residual vector quantization (RVQ) (Guo et al. 2024) with L+1L{+}1 levels:
@@
where $\text{sg}[\cdot]$ denotes the stop-gradient operation, and $\gamma$ controls the strength of the embedding loss.
@@
We design our 2D retrieval-augmented masked transformer to generate the base-layer motion tokens $\mathbf{y}^{0}\in\mathbb{R}^{T\times J}$, conditioned on the text embedding $t$, retrieved text feature $R_{t}$, and retrieved motion feature $R_{m}$, all of which are fused using SSTA. To train the model, we randomly mask a subset of tokens in $\mathbf{y}^{0}$, replacing them with a $\mathrm{MASK}]$ token to obtain a corrupted sequence $\mathbf{y}^{0}_{\text{msk}}$. We perform a 2D Mask strategy (Yuan et al. 2024) by first randomly masking along the temporal dimension and then randomly masking along the spatial dimension on those unmasked frames. The model is then trained to reconstruct the original tokens by minimizing the negative log-likelihood:
@@
During training, unconditional sampling is applied with a probability of $10\%$ to enable guidance-free learning. At inference, R-CFG is applied to the final projection layer prior to softmax.